OpenStudy (anonymous):

When you see a traffic light turn red, you apply the brakes until you come to a stop. If your initial speed was 12 m/s, and you were heading due west, what was your average velocity during breaking? Assume constant deceleration.

OpenStudy (anonymous):

Using the equations of linear motion under constant acceleration from the source below (vAV = average velocity): vAV = (v0 + v)/2 = (12 m/s + 0 m/s)/2 = 6 m/s West.

OpenStudy (anonymous):

OMG I feel retarded now! Thank you!

OpenStudy (anonymous):

No problem! :D